OpenStudy (anonymous):

what is the remainder of the function f(x)=x^3+x^2+x+2 divided by x- 4. show work. PLEASE HELP ME!!!!!!! ASAP

jimthompson5910 (jim_thompson5910):

you could use long division or synthetic division

jimthompson5910 (jim_thompson5910):

but a shortcut is to use the remainder theorem

OpenStudy (anonymous):

CAN YOU SHOW ME PLEASE

jimthompson5910 (jim_thompson5910):

which says when you divide a polynomial f(x) by x-c, the remainder is f(c)

jimthompson5910 (jim_thompson5910):

so what this means is that you just plug in x = 4 into f(x)

jimthompson5910 (jim_thompson5910):

f(x)=x^3+x^2+x+2 f(4)=(4)^3+(4)^2+4+2 f(4) = ???

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

86

jimthompson5910 (jim_thompson5910):

that's your remainder

OpenStudy (anonymous):

thank you this helped me out a lot

jimthompson5910 (jim_thompson5910):

you're welcome